# Data-driven model order reduction for wave propagation in materials with nonlinearities or damage

### v1

@@ Abstract
In this work, we consider wave propagation in materials characterized by nonlinear properties or damage. To accelerate the simulations of the resulting high-dimensional problems, we apply model order reduction methods. Depending on the knowledge of the underlying equations and the availability of their discrete operators, intrusive methods (here projection-based approaches based on proper orthogonal decomposition (POD)) or non-instrusive methods (here data-driven approaches including dynamic mode decomposition (DMD) and operator inference (OpInf)) can be used. We recall the theoretical foundations of the methods and apply them to the problem of wave propagation. In three different numerical examples, we evaluate the performance of the reduction techniques.
@@
We apply the methods proper orthogonal decomposition (POD), dynamic mode decomposition (DMD), multiresolution DMD (mrDMD) and operator inference (OpInf) to different problems of guided ultrasonic wave (GUW) propagation which are the classical wave equation, a mechanical system resulting from GUW propagation and a nonlinear hyperelastic model. The operator inference approach is modified by two aspects: we introduce a scaling approach for the snapshot data (see Section 3.5) and propose an unconstrained operator inference formulation which preserves structure and allows the employment of first order optimizers, see (36). Further, we continue and extend our study on the nonlinear aluminium model started in [GH24].
@@ 1 Introduction
Partial differential equations (PDEs) are fundamental tools in science and engineering, as they describe a wide range of physical phenomena derived from conservation laws such as mass and momentum [quarteroni2009numerical]. Their numerical solution typically relies on discretization techniques, most prominently the finite element method (FEM), which converts the PDE into a high-dimensional system of ordinary differential equations (ODEs) [zienkiewicz2005finite].
@@
In structural mechanics, the governing equation of motion takes the form of a PDE derived from the balance of momentum [hughes2012finite]. This PDE depends on parameters that influence the structural response, such as material properties, geometry, or boundary and initial conditions. In this work, we focus on parameters associated with material degradation, often referred to as damage parameters. The motivation stems from structural health monitoring (SHM), where the identification and localization of damage is essential to prevent catastrophic failures and to reduce maintenance costs [Farrar2006].
@@
Solving parameterized PDE problems for multiple parameter values is often computationally prohibitive. The FEM discretization typically yields systems of very large dimension, leading to high computational demands in terms of both time and memory. Repeated simulations, as required in real-time control, uncertainty quantification, inverse problems, or optimization, can therefore become infeasible. To address this issue, the scientific community has developed a wide range of model order reduction (MOR) techniques over the last decades [hesthaven2015certified, Bader2016, prud2002reliable, Barrault2004, Grepl2007, Benner2015, quarteroni2014reduced, quarteroniRB2016]. These methods aim to construct low-dimensional surrogate models that retain the essential dynamics of the high-fidelity system while significantly reducing computational costs. MOR has proven particularly effective for problems in computational mechanics, encompassing both structural and fluid dynamics applications [Rozza2008, Guo2018, Kunisch2002, Carlberg2013, Ballarin2016].
@@
In this article, we focus on the problem of wave propagation in damaged materials within the context of SHM [Farrar2006]. Non-destructive evaluation (NDE) techniques are commonly employed to detect and localize defects in structures. The objective of SHM is to leverage NDE methods to continuously monitor structural integrity in an automated manner, using appropriately placed sensors. Among the various approaches, a widely used technique is guided ultrasonic waves (GUW) [amafabia2017review]. In this method, guided waves are propagated through the structure, and their interaction with damage alters the signals recorded by the sensors [lammering2017lamb, su2006guided, guy2003guided, Fulgence2024, FulgenceNikiema2023]. The objective is to process these measured signals using numerical methods to characterize the presence, location, and severity of damage. We consider in this work two main models which are a linear wave equation (Section 2.1) and a linear mechanical system resulting from GUW propagation in materials with damage (Section 2.2). In addition, we consider wave propagation in nonlinear elastic material without defects (Section 2.4) as a first step to later include damage in that setting. The type of damage considered in this work is always a material degradation damage. For spatial discretization, we use the finite element method for both the wave equation and the equation of motion. In the case of the GUW simulations, the discretization in space and time is fine due to the high frequency of the simulation. This results in a high number of degrees of freedom of the discretized model raising the computational cost needed to perform the simulation. This motivates the use of model order reduction techniques in order to mitigate the computational costs associated with running the GUW simulations such that the reduced model can later be used to estimate the damage parameters solving an inverse problem.
@@
Model order reduction techniques can broadly be classified into intrusive and non-intrusive approaches. Intrusive methods require full knowledge of the governing equations and direct access to the discretized system matrices, while non-intrusive methods construct reduced models from solution data (and, in some cases, derivative data) without explicit access to the full order operators. In this work, we examine exemplary approaches from both categories in the context of the wave equation and subsequently focus on non-intrusive techniques for the mechanical system arising from the GUW simulations. Different non-intrusive methods are compared in terms of accuracy, computational efficiency, and additional performance metrics.
@@
Intrusive approaches in reduced order modeling include the well-known greedy reduced basis method [hesthaven2015certified], the proper orthogonal decomposition (POD) [volkwein2013proper, Bergmann2009516, Baiges2014189, Burkardt2006, Ballarin2016, Grle2017], and the proper generalized decomposition (PGD) [Dumon20111387, Chinesta2011], among others. These methods require a projection of the full order equations onto suitable reduced basis spaces, and are therefore also referred to as projection-based reduced order models.
@@
In contrast, non-intrusive methods can be broadly classified into two categories. The first relies solely on solution data to construct the reduced order model, while the second may require additional knowledge about the underlying system, such as the governing equations, forcing terms, or initial and boundary conditions. A representative example of the first category is the dynamic mode decomposition (DMD) method [Schmid2010], which is a purely data-driven technique that seeks the best linear operator to advance the discrete solution in time. Other fully data-driven approaches employ regression techniques or artificial neural networks (ANNs) to approximate the mapping between time/parameter inputs and the reduced basis projection coefficients [Loiseau2018, Guo2018, Hesthaven2018, Guo2019]. Non-intrusive approaches of the second category may require, in addition to data, some prior information about the structure of the system. A prominent example is the operator inference (OpInf) framework, which approximates the governing operators that generate the observed solution data. Originally introduced in [Peherstorfer2016Opinf], the approach has since been extended in various directions [Filanova2023, Benner2020, Qian2020, Yldz2021, SHARMA2024116865, Uy2023, Freitag2025].
@@
Finally, we also highlight hybrid reduced order models, which leverage both physical equations and solution data. Examples include calibration-based methods [noack2005, GALLETTI2004, Couplet2005], data-driven filtering and correction techniques [Xie2018, Ivagnes2023], POD with interpolation [Demo2019, Nguyen2022, Hijazi2020, Hijazi2020JCP], and the use of physics-informed neural networks (PINNs) at the reduced level [Chen2021, Hijazi2023].
@@ 2.1 Wave equation
The first model we consider is based on the classical wave equation, in which damage is incorporated via parameterization. Following [NandaLorenzWave], damage is characterized by its spatial coordinates and by a local reduction of the wave propagation speed. Here, we study a modification of this model as proposed in [Bur2024] by additionally introducing the damage size and considering a smoothing modification of the initial condition and the damage parameterization.
@@
Here, $u$ represents the displacement, $c$ is the wave speed coefficient and $f$ is the external forcing term. The initial conditions for the solution and the first time derivative at $t=t_{0}$, respectively, are given by $u_{0}$ and $u_{1}$, which specify the initial displacement and its first partial temporal derivative, respectively. On the boundary $\partial\Omega$, the displacement is prescribed via Dirichlet boundary conditions given through the function $g$. Details on the choices for the data used in the numerical simulations are provided in Section 4.1. For the numerical solution of (1), we use piecewise linear finite elements for discretization in space and the Crank-Nicolson scheme for discretization in time, see Section 2.4 for more details on this.
@@ 2.2 Mechanical system resulting from GUW propagation
The equation of motion is discretized using the finite element method, employing quadrilateral plane strain elements and quadratic Lagrange shape functions for spatial discretization. The time discretization is conducted within the commercial software of COMSOL Multiphysics®. In the numerical experiments, we consider the linear two-dimensional model of fiber metal laminate (FML) with damage considered in [BellamMuralidhar2021, BellamMuralidhar2023, Abschlussband2025], see Section 4.2 for more details.
@@ 2.3 Nonlinear hyperelastic model
The third model considered in this work involves hyperelastic material with nonlinearity, see e.g. [Belytschko2000]. For this purpose we give a brief introduction of how the nonlineariy is entering the mathematical formulation of the problem.
@@
If we assume hyperelastic material behavior, these materials are characterized by the existence of a strain energy density function $\Psi$ as the potential of the stresses [Belytschko2000, Hahn1985]. Based on this, the second Piola-Kirchoff stress tensor is given as:
@@
where $\lambda$ is the first Lamé parameter and $\mu$ is the shear modulus or the second Lamé parameter [Pence2014]. Then the resulting stress tensor is given by:
@@ 2.4 Discrete second order mathematical model
By $t$ we denote time and $[0,T]$ is the time interval of interest. The vector $\mathbf{u}\in\mathbb{R}^{N}$ collects the nodal displacements [^1] Note that we use the notation $\mathbf{u}$ to refer to both the continuous displacement field in (2) and its discrete counterpart in (9); the intended interpretation is clear from the context.. The matrices $\mathbf{M}\in\mathbb{R}^{N\times N}$ and $\mathbf{K}\in\mathbb{R}^{N\times N}$ are the mass and stiffness matrices, respectively, $\mathbf{f}\in\mathbb{R}^{N}$ is the external force vector, $\mathbf{g}$ is the nonlinear part when nonlinearity is present and $N$ is the number of degrees of freedom. Often, the right hand side $\mathbf{f}$ in (9) is expressed in terms of input signals $\mathbf{z}(t)\in\mathbb{R}^{N_{I}}$, with $N_{I}$ input signals, and an input-mapping matrix $\mathbf{B}\in\mathbb{R}^{N\times N_{I}}$ as follows:
@@
In our work, we are interested in studying how solutions of (9), (10) and (11) respond to variations in input parameters $\bm{\theta}\in\mathcal{D}\subset\mathbb{R}^{N_{p}}$, which could comprise the wave speed coefficient appearing in (1), the initial conditions, or material constants embedded in the equation of motion (2), which governs the virtual displacement in an elastic material—for example, the Young’s modulus of elasticity. In particular, by choosing the parameter vector $\bm{\theta}$, our model can represent a damage state in the material. We rewrite the systems in (9) and (10), respectively, to make the $\bm{\theta}$-dependency explicit:
$\mathbf{M}\ddot{\mathbf{u}}(\bm{\theta},t)+\mathbf{K}(\bm{\theta})\mathbf{u}(\bm{\theta},t)=\mathbf{f}(t),\quad\textrm{for}\quad\left(\bm{\theta},t\right)\in\mathcal{D}\times[0,T]$ (12)
and
$\mathbf{M}\ddot{\mathbf{u}}(\bm{\theta},t)+\mathbf{g}(\mathbf{u}(\bm{\theta},t))=\mathbf{f}(t),\quad\textrm{for}\quad\left(\bm{\theta},t\right)\in\mathcal{D}\times[0,T].$ (13)
Note that we assume the mass matrix $\mathbf{M}$ to be parameter independent while the stiffness matrix $\mathbf{K}$ varies with the parameter $\bm{\theta}$. Consequently, the displacement field $\mathbf{u}$ depends also on the parameter $\bm{\theta}$. This assumption comes from the problems of interest which will be addressed in the numerical results, see Section 4. For the rest of the paper, we treat the parameter $\bm{\theta}$ as fixed in each run. In future work, we intend to develop parameter dependent reduced order models for (12) and (13), see Section 5.
@@
$\mathbf{U}$ $=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}(\bm{\theta},t_{1})&\cdots&\mathbf{u}(\bm{\theta},t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M},\quad\mathbf{Z}=\begin{bmatrix}|&\cdots&|\\ \mathbf{z}(t_{1})&\cdots&\mathbf{z}(t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N_{I}\times M}.$ (14)
@@ 3 Model order reduction techniques
In this work, we consider a common challenge that often appears in computational sciences: time series solution data of the state variable are generated by a numerical solver to which the user may not have full access. Such a situation motivates the search for an approximation of the mathematical system which produced the numerical data. Operator inference (OpInf, see Section 3.2) utilizes knowledge of the structure of the algebraic form in order to infer reduced operators directly from solution snapshots. In case we do not have any knowledge of the structure of the underlying dynamical system, dynamic mode decomposition (DMD, see Section 3.6) can be considered which is entirely data-driven and extracts coherent spatio-temporal modes and associated dynamics. Both operator inference and DMD are non-intrusive approaches as they do not require the full order system matrices. In contrast, classical intrusive MOR methods like POD-Galerkin (see Section 3.3) demand access to the full order matrices before reduction. In the following, we briefly recall these three approaches and then demonstrate their application to the second order models (9), (10) and (11).
@@ 3.1 Snapshot data
Let us introduce the notation for the snapshot data to which we refer in the subsequent chapters. At given equidistant time instants $0=t_{1}<\dots<t_{M}=T$ the data is gathered which consists of the input signals $\mathbf{z}(t)$ and the nodal displacements of the mechanical system collected in the vector $\mathbf{u}(\bm{\theta},t)$ at the given time points $t=t_{i}$ for $i=1,\dots,M$ for a fixed parameter $\bm{\theta}$. The snapshot matrices of the displacement and input signals are formed as follows:
$\mathbf{U}$ $=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}(\bm{\theta},t_{1})&\cdots&\mathbf{u}(\bm{\theta},t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M},\quad\mathbf{Z}=\begin{bmatrix}|&\cdots&|\\ \mathbf{z}(t_{1})&\cdots&\mathbf{z}(t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N_{I}\times M}.$ (15)
The basic idea of operator inference is to recover a numerical model directly from snapshot data by approximating the system operators. Usually, the structure of the system that resulted in the numerical data is known to a certain extent. This partial knowledge is used to formulate a mathematical problem for the approximation of the system operators. Since the estimation of the full order operators would be computationally immense, one directly estimates reduced system operators which are of much smaller dimension. In [Peherstorfer2016Opinf], this approach has been proposed and is termed as operator inference. The operator inference approach has also been extended to the non-linear case and to specific physical problems, see, e.g. [Benner2020, Qian2020, Yldz2021]. An operator inference method tailored for second order mechanical systems has been proposed in [Filanova2023], which we closely follow here to apply it to our mechanical systems of interest.
In the operator inference approach, derivative information of the solution data is required. In many occasions, the numerical solver may provide this data. If that is not the case, then one may approximate the derivatives using finite difference schemes, e.g. using an eighth order finite difference scheme, see [Sharma2024]. The second derivatives of the displacement are assembled in the following snapshot matrix:
$\ddot{\mathbf{U}}$ $=\begin{bmatrix}|&\cdots&|\\ \ddot{\mathbf{u}}(\bm{\theta},t_{1})&\cdots&\ddot{\mathbf{u}}(\bm{\theta},t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M}.$ (16)
$\mathbf{\widehat{U}}=\mathbf{\Phi}^{T}\mathbf{U},\quad\ddot{\mathbf{\widehat{U}}}=\mathbf{\Phi}^{T}\ddot{\mathbf{U}},$ (17)
where $\mathbf{\Phi}$ is the matrix of POD modes computed using e.g. a singular value decomposition (SVD) of $\mathbf{U}$, see subsection 3.3.

### v2

# Data-driven model order reduction for wave propagation in materials with damage and nonlinearities

@@ Abstract
In this work, we consider wave propagation in materials characterized by nonlinear properties and damage. To accelerate the simulations of the resulting high-dimensional problems, we apply model order reduction methods. Depending on the knowledge of the underlying equations and the availability of their discrete operators, intrusive methods (here projection-based approaches based on proper orthogonal decomposition (POD)) or non-instrusive methods (here data-driven approaches including dynamic mode decomposition (DMD) and operator inference (OpInf)) can be used. We recall the theoretical foundations of the methods and apply them to the problem of wave propagation. In three different numerical examples, we evaluate the performance of the reduction techniques.
@@
We apply the methods proper orthogonal decomposition (POD), dynamic mode decomposition (DMD), multiresolution DMD (mrDMD) and operator inference (OpInf) to different problems of guided ultrasonic wave (GUW) propagation which are the classical wave equation, a mechanical system resulting from GUW propagation and a nonlinear hyperelastic model. The operator inference approach is modified by following aspects: we propose an unconstrained operator inference formulation which preserves structure and allows the employment of first order optimizers, see (39). We also present a time discrete operator inference approach for second order mechanical systems. In addition, the re-projection method which recovers the intrusive reduced operators has been extended for second order mechanical systems. Further, we continue and extend our study on the nonlinear aluminium model started in [29], a damage in the nonlinear model is introduced in this work.
@@ 1 Introduction
Partial differential equations (PDEs) are fundamental tools in science and engineering, as they describe a wide range of physical phenomena derived from conservation laws such as mass and momentum [71]. Their numerical solution typically relies on discretization techniques, most prominently the finite element method (FEM), which converts the PDE into a high-dimensional system of ordinary differential equations (ODEs) [85].
@@
In structural mechanics, the governing equation of motion takes the form of a PDE derived from the balance of momentum [45]. This PDE depends on parameters that influence the structural response, such as material properties, geometry, or boundary and initial conditions. In this work, we focus on parameters associated with material degradation, often referred to as damage parameters. The motivation stems from structural health monitoring (SHM), where the identification and localization of damage is essential to prevent catastrophic failures and to reduce maintenance costs [25].
@@
Solving parameterized PDE problems for multiple parameter values is often computationally prohibitive. The FEM discretization typically yields systems of very large dimension, leading to high computational demands in terms of both time and memory. Repeated simulations, as required in real-time control, uncertainty quantification, inverse problems, or optimization, can therefore become infeasible. To address this issue, the scientific community has developed a wide range of model order reduction (MOR) techniques over the last decades [39, 3, 67, 6, 32, 10, 70, 69]. These methods aim to construct low-dimensional surrogate models that retain the essential dynamics of the high-fidelity system while significantly reducing computational costs. MOR has proven particularly effective for problems in computational mechanics, encompassing both structural and fluid dynamics applications [75, 33, 49, 16, 5].
@@
In this article, we focus on the problem of wave propagation in damaged materials within the context of SHM [25]. Non-destructive evaluation (NDE) techniques are commonly employed to detect and localize defects in structures. The objective of SHM is to leverage NDE methods to continuously monitor structural integrity in an automated manner, using appropriately placed sensors. Among the various approaches, a widely used technique is guided ultrasonic waves (GUW) [2]. In this method, guided waves are propagated through the structure, and their interaction with damage alters the signals recorded by the sensors [52, 79, 35, 59, 58]. The objective is to process these measured signals using numerical methods to characterize the presence, location, and severity of damage. We consider in this work two main models which are a linear wave equation (Section 2.1) and a linear mechanical system resulting from GUW propagation in materials with damage (Section 2.2). In addition, we consider wave propagation in nonlinear elastic material without defects (Section 2.4) as a first step to later include damage in that setting. The type of damage considered in this work is always a material degradation damage. For spatial discretization, we use the finite element method for both the wave equation and the equation of motion. In the case of the GUW simulations, the discretization in space and time is fine due to the high frequency of the simulation. This results in a high number of degrees of freedom of the discretized model raising the computational cost needed to perform the simulation. This motivates the use of model order reduction techniques in order to mitigate the computational costs associated with running the GUW simulations such that the reduced model can later be used to estimate the damage parameters solving an inverse problem.
@@
Model order reduction techniques can broadly be classified into intrusive and non-intrusive approaches. Intrusive methods require full knowledge of the governing equations and direct access to the discretized system matrices, while non-intrusive methods construct reduced models from solution data (and, in some cases, derivative data) without explicit access to the full order operators. In this work, we examine exemplary approaches from both categories in the context of the wave equation and subsequently focus on non-intrusive techniques for the mechanical system arising from the GUW simulations due to the restricted access to the system’s operators using a commercial software. Different non-intrusive methods are compared in terms of accuracy, computational efficiency, and additional performance metrics.
@@
Intrusive approaches in reduced order modeling include the well-known greedy reduced basis method [39], the proper orthogonal decomposition (POD) [81, 11, 4, 14, 5, 30], and the proper generalized decomposition (PGD) [23, 19], among others. These methods require a projection of the full order equations onto suitable reduced basis spaces, and are therefore also referred to as projection-based reduced order models.
@@
In contrast, non-intrusive methods can be broadly classified into two categories. The first relies solely on solution data to construct the reduced order model, while the second may require additional knowledge about the underlying system, such as the governing equations, forcing terms, or initial and boundary conditions. A representative example of the first category is the dynamic mode decomposition (DMD) method [76], which is a purely data-driven technique that seeks the best linear operator to advance the discrete solution in time. Other fully data-driven approaches employ regression techniques or artificial neural networks (ANNs) to approximate the mapping between time/parameter inputs and the reduced basis projection coefficients [53, 33, 38, 34]. Non-intrusive approaches of the second category may require, in addition to data, some prior information about the structure of the system. A prominent example is the operator inference (OpInf) framework, which approximates the intrusive reduced operators allowing for the construction of a low rank system that can be integrated efficiently to perform prediction tasks both in time and parameter. Originally introduced in [63], the approach has since been extended in various directions [26, 9, 68, 84, 78, 80, 27].
It is well-known that hyperbolic problems and systems involving transport phenomena are challenging for model order reduction due to slow decaying Kolmogorov n-width, i.e. no choice of an n-dimensional linear subspace can lead to a faster error decay than the given width, see e.g. [61, 12, 31]. In order to overcome this barrier, nonlinear model reduction methods can be employed which are not restricted to fixed linear subspaces, see e.g. [65, 40].
@@
Finally, we also highlight hybrid reduced order models, which leverage both physical equations and solution data. Examples include calibration-based methods [60, 28, 21], data-driven filtering and correction techniques [83, 46], POD with interpolation [22, 57, 41, 44], and the use of physics-informed neural networks (PINNs) at the reduced level [18, 43].
@@ 2.1 Wave equation
The first model we consider is based on the classical wave equation, in which damage is incorporated via parameterization. Following [55], damage is characterized by its spatial coordinates and by a local reduction of the wave propagation speed. Here, we study a modification of this model as proposed in [15] by additionally introducing the damage size and considering a smoothing modification of the initial condition and the damage parameterization.
@@
Here, $u$ represents the displacement, $c$ is the wave speed coefficient and $f$ is the external forcing term. The initial conditions for the solution and the first time derivative at $t=t_{0}$, respectively, are given by $u_{0}$ and $u_{1}$. On the boundary $\partial\Omega$, the displacement is prescribed via Dirichlet boundary conditions. For the numerical solution of (1), we use piecewise linear finite elements for discretization in space, resulting in a system of ordinary differential equations, see Section 2.4. For the discretization in time, we use the Crank-Nicolson scheme. Details on the choices for the data used in the numerical simulations are provided in Section 4.1
@@ 2.2 Mechanical system resulting from GUW propagation
The equation of motion is discretized using the finite element method, employing quadrilateral plane strain elements and quadratic Lagrange shape functions for spatial discretization. The time discretization is conducted within the commercial software of COMSOL Multiphysics®. In the numerical experiments, we consider the linear two-dimensional model of fiber metal laminate (FML) with damage considered in [56, 54, 42], see Section 4.2 for more details.
@@ 2.3 Nonlinear hyperelastic model
The third model considered in this work involves hyperelastic material with nonlinearity, see e.g. [7]. For this purpose we give a brief introduction of how the nonlinearity enters the mathematical formulation of the problem.
@@
If we assume hyperelastic material behavior, these materials are characterized by the existence of a strain energy density function $\Psi$ as the potential of the stresses [7, 37]. Based on this, the second Piola-Kirchoff stress tensor is given as:
@@
where $\lambda$ is the first Lamé parameter and $\mu$ is the shear modulus or the second Lamé parameter [66]. Then, the resulting stress tensor is given by:
@@ 2.4 Discrete second order mathematical model
By $t$ we denote time and $[0,T]$ is the time interval of interest. The vector $\mathbf{u}\in\mathbb{R}^{N}$ collects the nodal displacements [^1] Note that we use the notation $\mathbf{u}$ to refer to both the continuous displacement field in (2) and its discrete counterpart in (9); the intended interpretation is clear from the context.. The matrices $\mathbf{M}\in\mathbb{R}^{N\times N}$ and $\mathbf{K}\in\mathbb{R}^{N\times N}$ are the mass and stiffness matrices, respectively, $\mathbf{f}(t)\in\mathbb{R}^{N}$ is the external force vector, $\mathbf{g}$ is the nonlinear part and $N$ is the number of degrees of freedom. Often, the right hand side $\mathbf{f}$ in (9) is expressed in terms of input signals $\mathbf{z}(t)\in\mathbb{R}^{N_{I}}$, with $N_{I}$ input signals, and an input-mapping matrix $\mathbf{B}\in\mathbb{R}^{N\times N_{I}}$ as follows:
@@ 3 Model order reduction techniques
In this work, we consider a common challenge that often appears in computational sciences: time series solution data of the state variable are generated by a numerical solver to which the user may not have full access. Such a situation motivates the search for an approximation of the mathematical system which produced the numerical data. Operator inference (OpInf, see Section 3.3 and Sections 3.4, 3.5, 3.6, 3.7) utilizes knowledge of the structure of the underlying problem in order to infer reduced operators directly from solution snapshots. In case we do not have any knowledge of the structure of the underlying dynamical system, dynamic mode decomposition (DMD, see Section 3.8) can be considered which is entirely data-driven and extracts coherent spatio-temporal modes and associated dynamics. Both operator inference and DMD are non-intrusive approaches as they do not require the full order system matrices. In contrast, classical intrusive model order reduction (MOR) methods like proper orthogonal decomposition (POD) based Galerkin approaches (see Section 3.2) demand access to the full order matrices before reduction. In the following, we briefly recall these three approaches and then demonstrate their application to the second order models (9), (10) and (11).
@@ 3.1 Snapshot data
Let us introduce the notation for the snapshot data to which we refer in the subsequent chapters. At given equidistant time instants $0=t_{1}<\dots<t_{M}=T$ the data is gathered which consists of the input signals $\mathbf{z}(t)$ and the nodal displacements of the mechanical system collected in the vector $\mathbf{u}(t)$ at the given time points $t=t_{i}$ for $i=1,\dots,M$. The snapshot matrices of the displacement and input signals are formed as follows:
$\mathbf{U}$ $=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}(t_{1})&\cdots&\mathbf{u}(t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M},\quad\mathbf{Z}=\begin{bmatrix}|&\cdots&|\\ \mathbf{z}(t_{1})&\cdots&\mathbf{z}(t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N_{I}\times M}.$ (12)